Complete the pattern 43,37,31,25,_,_
Answer by mark_henrhiz(40)

this is not a mixture problem
in these kinds of number series, always find a pattern
the pattern with this one, is you subtract 6 from the preceding term to get the term after it.
so,
43 - 6 = 37
37 - 6 = 31
31 - 6 = 25
25 - 6 = 19
19 - 6 = 13

19, 13